Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $130 per unit.
Answer #1

1.

Selling price per unit = $130

Variable cost per unit = $91

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 130-91

= $39

Contribution margin ratio = Contribution margin per unit / Selling price pr unit

= 39/130

= 30%

Break even point (in units) = Fixed costs / Contribution margin per unit

= 179,400/39

= 4,600 units

Break even point (in $) = Fixed costs / Contribution margin ratio

= 179,400/30%

= $598,000

Break-Even Point
Number of stoves 4,600
Total sales dollars $598,000

2.

When variable expense per stoves increase as a percentage of the selling price, it will result in higher break even point since due to increase in variable expenses, lower contribution margin per unit will be available, leading to higher break even point.

3.

Reduction in selling price per unit = 10%

= 130 x 10%

= $13

Proposed selling price per unit = 130-13

= $117

Increase in sales units = 25%

= 17,000 x 25%

= 4,250 units

Proposed sales units = 17,000+4,250

= 21,250 units

Outback Outfitters
Present Proposed
17,000 Stoves 21,250 Stoves
Total Per unit Total Per unit
Sales 2,210,000 130              2,486,250 117
Variable cost -1,547,000 91 -1,933,750 91
Contribution margin $663,000 39 $552,500 26
Fixed cost -179,400 -179,400
Operating Income $483,600 $373,100

4.

Number of stoves to be sold for target income = ( Fixed costs+ target income)/Contribution margin per unit

= (!79,400+78,000)/26

= 257,400/26

= 9,900 units

Kindly comment if you need further assistance.

Thanks‼!